Example 2 Find the Fourier Series coefficients for y(t) = sin? 20,1 + cos² ept =
Answer #1

2 @ YIE = sin? (zwot) + cost wat) using cos2o - Caso-sine Cos 20 = 1-251020 - and cosso = 200520 - . Cos 20 = 1+ cos20 5120 =

Hope it helps you out.

In case any doubt please feel free to ask.

I Will be happy to help you.

Thanks!!